OpenStudy (anonymous):

One more +1 riddle a train 100m long long moving on a straight level track passes a pole in just 5 s. Find: a) The Speed of the train b) the time it will take to cross a bridge of 500 m long.

OpenStudy (anonymous):

here comes lgba

OpenStudy (lgbasallote):

a) 100m/5s = 20 m/s b) 500/5 = 100 seconds

OpenStudy (lgbasallote):

oops lol b is wrong

OpenStudy (lgbasallote):

b) 500/20 = 25 seconds

OpenStudy (anonymous):

thanks dude to attend my question and your <tips hat>